# Few-shot Species Range Estimation

### v1

@@ 3.2.1 Additional Context Information
The design of FS-SINR is flexible, in that we can also provide additional context information to the model if it is available. For example, if there is additional text (e.g., a range description) or visual (e.g., images) information available for a novel species it can be added to the context, assuming such information was also available at training time for other species. This observation is inspired by recent work that also uses language derived information to improve range predictions (Sastry et al., 2023; Hamilton et al., 2024) and work that uses species images locations observations (Sastry et al., 2025). This additional information can provide a rich source of metadata encoding aspects of a species’ habitat preferences, even when there might only be a limited number of location observations available for it. We can represent the expanded contextual input tokens as $\{\bm{t}_{j},\bm{a}_{j},f_{\bm{\theta}}(\bm{c}_{1}),\dots,f_{\bm{\theta}}(\bm{c}_{k})\}$, where $\bm{t}_{j}$ denotes a fixed length text embedding from a large language model and $\bm{a}_{j}$ an image embedding obtained from a pretrained vision model for species $j$. Note, we train FS-SINR such that it can use arbitrary subsets of these input tokens during inference.
@@ 4.1 Implementation Details
Architecture. Our location encoders use the same fully connected neural network with residual connections as in Cole et al. (2023). Each of the context locations is processed by the same shared location encoder which is first pretrained as in SINR after which the multi-label classifier head is discarded. Importantly, this pretrained encoder is only trained on species from the training set, and does not observe any data from the evaluation species. The text embedding backbone is a frozen GritLM (Muennighoff et al., 2025) and the image embedding backbone is a frozen EVA-02 ViT (Fang et al., 2024) pretrained on the iNaturalist species classification dataset (Van Horn et al., 2021). Both backbones provide a fixed length embedding vector and we train two layer fully connected text and image encoders to transform these embeddings into their context tokens. FS-SINR’s Transformer contains four encoder layers and the parameters are updated jointly with the location, text, and image encoders and species decoder during training. In total, our model has 8.2M learnable parameters compared to 11.9M for SINR. We train with a batch size of 2,048 instances and randomly drop-out text/image or location tokens during training with a probability of 0.5 and 0.1 respectively to enhance robustness. See Section A.1 for more details.
@@
Figure 3: Few-shot results. Here we evaluate different models on the task of species range estimation on the IUCN (left) and S&T (right) datasets. On the x-axis we vary the number of context locations seen at inference time for the held-out evaluation species. The y-axis represents MAP, where higher values are better. The error bars display the standard deviation of three different runs. Our FS-SINR approach outperforms existing methods, especially in the very low-setting setting (i.e., $<$ five context locations). Note that LE-SINR and SINR need to be retrained during evaluation when more observations are provided.
@@
Data. We train FS-SINR using the presence-only dataset from Cole et al. (2023) which contains 35.5 million citizen science observations for 47,375 species from the iNaturalist platform iNaturalist (2025), along with 127 thousand text descriptions used in Hamilton et al. (2024) and 200 thousand images of the species which we obtain from iNaturalist (2025). During training, we supply FS-SINR with 20 context locations per training example, though we find that the model performance is very robust to the number of context locations provided during training. We evaluate models using the IUCN and S&T datasets also from Cole et al. (2023), which contain 2,418 and 535 expert and model-derived binary species range maps respectively. The IUCN dataset is more globally distributed and contains a larger variation in range size across species, while the S&T dataset only contains bird species that are primarily, but not always, found in North America and have a larger average range size. While not perfect, these datasets represent the currently best available evaluation data and contain large variety in terms of range sizes and locations. Importantly, unless otherwise stated, we hold out any species from the union of these two datasets from the training set so that species from the evaluation set are not observed during training. As a result, by default, FS-SINR is trained on data from 44,422 species. Performance is reported as mean average precision (MAP) for different numbers of input (i.e., context) locations.
@@
Baselines. Generating a species’ range from FS-SINR for a held-out species at inference time only requires a single forward pass through the model to obtain an embedding vector for the species. Current methods (e.g., LE-SINR or SINR) cannot be used in such a feed forward manner and need to be retrained for each species that was not observed at training time. To obtain an equivalent embedding for the SINR and LE-SINR baselines we train a per-species binary logistic regression classifier using any few-shot presence observations that are available, in addition to adding 10,000 uniformly random and 10,000 target (i.e., in locations where species are) pseudo-absences as in LE-SINR. The Prototype and Active SINR’s baselines do not require retraining. For fairness, we use the same presence observations across each method and the larger number of presences are supersets of the smaller ones. Implementation details of the baseline methods can be found in Section A.2.
@@ 4.2 Few-shot Evaluation
First we evaluate how effective different range estimation models are at few-shot range estimation. The goal for each model is to generate a plausible prediction for a previously unseen species’ range given limited location observations. Quantitative results are presented in Figure 3.
@@
The SINR baseline performs poorly in the low data regime, but as more data is added performance improves. As noted earlier, here a per-species embedding vector is learned using logistic regression using the provided presence locations and generated pseudo-absences. The recently introduced LE-SINR approach extends the basic SINR model to use text information (here range text), when available, at inference time. We can see that when any form of text data is available, LE-SINR outperforms SINR. We also compare to the species embedding combination method from Lange et al. (2023), denoted as Active SINR, and a Prototypical Network-style baseline (Snell et al., 2017). Like our FS-SINR approach, neither of these methods require retraining at inference time, but perform much worse.
In all instances, when the same metadata is available, FS-SINR outperforms existing methods. Furthermore, we also outperform SINR in the larger data regime (i.e., when 50 observations are available). In general we observe that image information is not as informative as text and does not help on average outside of the zero-shot case. This can be explained by the fact that a range text description provides much more context than an image of a previously unseen species. Importantly, unlike SINR and LE-SINR, FS-SINR does not need to be retrained at inference time. Instead, it can make predictions in a feed forward manner irrespective of the context data available. This is advantageous in interactive settings, whereby the model can compute the location embeddings for all query locations on earth once and then a user could experiment by adding different context information interactively.
We present qualitative results for three different species in Figure 7 where we visualize FS-SINR’s predictions as we change the number of context locations. Given only a single context location, the model does a sensible job at localizing the species on the earth. This supports the findings from Figure 3 where we observe strong performance even when only one context location is available. When more information is provided, the predicted range more closely resembles the expert-derived range shown in the first row. However, we do note that the model can still make mistakes in our low data setting, such as the erroneous predictions for the ‘Black and White Warbler’ in South America. In Figure 4 we illustrate some examples of how text information, when paired with limited context locations, can influence the model predictions. We observe dramatically different predicted ranges when the text prompt encourages the model to focus on different habitat types. We note that each of the predicted ranges are still consistent with the location of the single context location provided. Finally, in Figure 6 we compare FS-SINR range predictions to other approaches, namely SINR, LS-SINR, and Active SINR. We see that for this species FS-SINR more closely resembles the expert range when only three context locations are provided. Additional qualitative examples are provided in Appendix C.
@@ 4.3 Zero-shot Evaluation
In addition to being able to generate range predictions in the few-shot setting when limited location observations are provided, FS-SINR can also make predictions when no location information is provided, i.e., the zero-shot setting. These zero-shot results are presented in Table 1 for both the IUCN and S&T datasets.
Table 1: Zero-shot results. We report zero-shot performance where no location information is provided to each model, comparing to SINR (Cole et al., 2023) and LE-SINR (Hamilton et al., 2024). We denote additional metadata used by models as RT for ‘Range Text’, HT for ‘Habitat Text’, and ‘I’ for ‘Image’. TST represents ‘Test Species in Train’, indicating that a model uses location observations for the evaluation species at training time (e.g., SINR which provides an upper bound on performance), unlike other models where these species are excluded. TRT models are trained using taxonomic rank text as in Sastry et al. (2023), which are also provided with the full taxonomic description from ‘class’ to ‘species’ during evaluation. Results are presented as MAP, where higher is better.
@@
Figure 5: Range predictions using a single context image as input. We can condition FS-SINR on an arbitrary input image with no context locations or text, e.g., (left) a held-out species, (middle) a landmark, or (right) a landscape image.
@@
Figure 6: Qualitative comparison of Range Predictions for different methods. Predicted ranges from three context locations denoted as ‘◦’ and no additional metadata for the Black-naped Monarch. From left to right, FS-SINR (ours) with expert range inset, SINR, LE-SINR, and Active SINR. Please zoom in to see details.
@@
We report results for several variants of FS-SINR where different types of metadata data are used. As a baseline, we also present the performance of SINR (row 1) where the the evaluation species are part of its training set. We can also add data from these species to the training set of our approach which unsurprisingly boosts performance (i.e., row 3 vs. 9), though unlike SINR, FS-SINR does not have weights associated with individual species and so the impact of seeing evaluation species during training is fairly small. As a trivial baseline, we also report performance of FS-SINR (row 4) when no location or text metadata are provided, i.e., this is simply the output of the class token. As expected, this model performs poorly, but interestingly it seems to have learned some spatial prior that results in non-trivial predictions on S&T which contains bird species mostly concentrated in North America. We also compare to a version of FS-SINR (row 5) where we use taxonomic text as in LD-SDM (Sastry et al., 2023) (see Section B.7 for further details).
@@
In all instances, our FS-SINR approach outperforms LE-SINR, even when both models are provided with the same information at inference and training time (i.e., row 6 vs. 7 or row 8 vs. 9). Confirming observations from the LE-SINR paper, we see that range text is more informative than habitat text (i.e., row 7 vs. 9). Additionally, image information provides some non-trivial signal (i.e., row 4 vs. 10), but it is not as informative as text (i.e., row 9 vs. 10). As we can see in Figure 5, zero-shot image predictions can be sensible, but predicting an unobserved species’ range from a single input image is ill posed.
4.4 Ablations
@@
We provide additional ablation experiments for FS-SINR in Appendix B. We present results where we evaluate the impact of different input features, location encoders, and report performance using alternative evaluation metrics. There, we evaluate the impact of the amount of data used to train FS-SINR and pretrain the SINR location encoder we use. Finally, we also explore architectural modifications such as removing the final species decoder that operates on the output of the Transformer. We observe that FS-SINR is robust to these changes, justifying the design decision we made. We also provide a more ecologically relevant breakdown of results by reporting performance against range size, geographic location, and by taxonomic group.
@@
Refer to caption Refer to caption Refer to caption
@@ 5 Limitations
While FS-SINR exhibits impressive zero and few-shot performance, there are some limitations. First, given a set of input context locations FS-SINR is deterministic in that it will always generate the same output range map. In practice, in the few-shot regime, the same set of points could actually be representative of many different possible range maps. An obvious, and easy to implement, extension of our work is to introduce stochasticity into the model outputs, e.g., by treating class token output from the Transformer as a latent embedding for an additional sampling step. In Figure A16 we obverse that initializing FS-SINR with different random seeds during training results in diverse range predictions across the different models. We leave further exploration of this for future work. Second, at inference time, users may wish to provide example locations indicating where a specific species has not been found, i.e., confirmed absences. Currently our model is trained using presence-only data, but could be adapted to use absence information, if available, which could be denoted via a different embedding type vector which would be learned during training alongside our existing token type embeddings. However, obtaining reliable large-scale absence data for tens of thousands of species is a challenging task.
Finally, global-scale citizen science datasets like the one we use to train FS-SINR can contain large biases (Geldmann et al., 2016), e.g., location, temporal, or taxonomic biases, among others. We do not explicitly account for these biases during training, and thus we would caution the use of the predictions of our model in any applications that would use our range predictions in the context of biodiversity assessments. However, we note that we outperform existing and recent state-of-the-art range estimation methods, especially in the low observation data setting, and do not require any retraining at inference time.

### v2

@@ 3.2.1 Additional Context Information
The design of FS-SINR is flexible in that we can also provide additional context information to the model if it is available. For example, if there is additional text (e.g., a range description) or visual (i.e., images) information available for a novel species, it can be added to the context, assuming that such information was also available at training time for other species. This observation is inspired by recent work that also uses language-derived information to improve range predictions (Sastry et al., 2023; Hamilton et al., 2024) and work that uses species images and observations (Sastry et al., 2025). This additional information can provide a rich source of metadata encoding aspects of a species’ habitat preferences, even when there might only be a limited number of location observations available for it. We can represent the expanded contextual input tokens as $\{\bm{t}_{j},\bm{a}_{j},f_{\bm{\theta}}(\bm{c}_{1}),\dots,f_{\bm{\theta}}(\bm{c}_{k})\}$, where $\bm{t}_{j}$ denotes a fixed-length text embedding from a large language model and $\bm{a}_{j}$ an image embedding obtained from a pre-trained vision model for species $j$ – see Figure 2. Note that we train FS-SINR so that it can use arbitrary subsets, including none, of these input tokens during inference.
@@ 4.1 Implementation Details
Architecture. Our location encoders use the same fully connected neural network with residual connections as in Cole et al. (2023). Each of the context locations is processed by the same shared location encoder which is first pre-trained as in SINR after which the multi-label classifier head is discarded. Importantly, this pre-trained encoder is only trained on species from the training set, and does not observe any data from the evaluation species during training. The text embedding backbone is a frozen GritLM (Muennighoff et al., 2025) and the default image embedding backbone is a frozen EVA-02 ViT (Fang et al., 2024) pre-trained on the iNaturalist species image classification dataset (Van Horn et al., 2021). Both backbones provide a fixed length embedding vector, and we train two-layer fully connected text and image encoders to transform these embeddings into their context tokens. FS-SINR’s Transformer contains four encoder layers and the parameters are updated jointly with the location, text, and image encoders and species decoder during training. In total, FS-SINR has 8.2M learnable parameters compared to 11.9M for SINR. This reduction is due to the fact that we do not have to learn a per-species embedding vector as in SINR. We train with a batch size of 2,048 instances and randomly drop-out text/image or location tokens during training with a probability of 0.5 and 0.1 respectively to enhance robustness. See Section C.1 for more details. Code for FS-SINR is available at: https://github.com/Chris-lange/fs-sinr
@@
Figure 3: Few-shot results. Here we evaluate different models on the task of species range estimation on the IUCN (left) and S&T (right) datasets. On the x-axis we vary the number of context locations seen at inference time for the held-out evaluation species. The y-axis represents Mean Average Precision (MAP), where higher values are better. The error bars display the standard deviation of three different runs. Our FS-SINR approach outperforms existing methods, especially in the very low-data setting (i.e., $<$ five context locations). Note that LE-SINR and SINR need to be retrained during evaluation when more observations are provided. Tables A3 and A4 report expanded results including larger numbers of context locations.
Data. We train FS-SINR on the presence-only dataset from Cole et al. (2023), which comprises 35.5 million citizen-science records—each annotated with latitude, longitude, and species label—for 47,375 diverse species including plants, fungi, and animals from the iNaturalist platform (iNaturalist, 2025). We also leverage 127 thousand text descriptions of these species used in  Hamilton et al. (2024) and 200 thousand images obtained from iNaturalist (iNaturalist, 2025). The text provided during training is composed of sections of Wikipedia (Wikipedia, 2025) articles of the target species. During training, we supply FS-SINR with 20 context locations per training example, although we find that model performance is robust to changes in the number of context locations provided during training.
@@
We evaluate models using the IUCN and S&T datasets also from Cole et al. (2023), which contain expert and model-derived range maps for 2,418 and 535 different species, respectively. The IUCN dataset is more globally distributed and contains a larger variation in range size and more diverse animal species, while the S&T dataset only contains bird species that are found primarily, but not always, in North America and have a larger average range size. We follow the same preprocessing steps for these datasets as in Cole et al. (2023). While not perfect, these datasets represent the best evaluation data currently available and contain large variety in terms of range sizes and locations. The text used during evaluation consists of pre-trained large language model generated summaries of the range or habitat of the target species as used in Hamilton et al. (2024). Importantly, we hold out any species from the union of these two datasets from the training set so that species from the evaluation set are not observed during training. As a result, by default, FS-SINR is trained on data from 44,422 species. Performance is reported as mean average precision (MAP) for different numbers of input (i.e., context) locations.
@@
Baselines. Generating a species’ range from FS-SINR for a held-out species at inference time only requires a single forward pass through the model to obtain an embedding vector for the species. Current methods (e.g., LE-SINR or SINR) cannot be used in such a feedforward manner and need to be retrained for each species that was not observed at training time. To obtain an equivalent embedding for the SINR and LE-SINR baselines we train a per-species binary logistic regression classifier using any few-shot presence observations that are available, in addition to adding 10,000 uniformly random and 10,000 target (i.e., in locations where species are) pseudo-absences as in LE-SINR. We also compare to the species embedding combination method from Lange et al. (2023) and a Prototypical Network-style baseline (Snell et al., 2017), denoted Active SINR and Prototype SINR, respectively. These baselines do not require retraining. For fairness, we use the same presence observations across each method, and the larger number of presences are supersets of the smaller ones. Implementation details of the baseline methods can be found in Section C.2.
@@ 4.2 Few-shot Evaluation
First, we evaluate how effective different range estimation models are at few-shot range estimation. The goal for each model is to generate a plausible prediction for a previously unseen species’ range given limited location observations. Quantitative results are presented in Figure 3, and additional results can be found in Appendix A.
The SINR baseline performs poorly in the low-data regime, but as more data is added performance improves. As noted earlier, here a per-species embedding vector is learned using logistic regression using the provided presence locations and generated pseudo-absences. The recently introduced LE-SINR approach extends the basic SINR model to use text information (here range text), when available, at inference time. LE-SINR tends to outperform SINR, particularly when text data is available. Like our FS-SINR approach, neither the Active SINR or Prototype SINR baselines require retraining at inference time, but perform much worse than FS-SINR.
In all instances, when the same metadata is available, FS-SINR outperforms existing methods. Furthermore, we also outperform SINR in the larger data regime (i.e., when 50 observations are available). In general, we observe that image information is not as informative as text and does not help on average outside of the zero-shot case. A range text description provides much more context than an image of a previously unseen species. However, range estimates for some species benefit significantly from using images, but other species actually see no change or decreased performance. Making use of images to improve zero-shot range estimates is rewarded during training, but this can harm performance for some species at inference time by biasing the ranges produced toward the rough estimates made using images, which can provide limited information. Importantly, unlike SINR and LE-SINR, FS-SINR does not need to be retrained at inference time. Instead, it can make predictions in a feedforward manner irrespective of the context data available. This is advantageous in interactive settings, whereby the model can compute the location embeddings for all query locations on earth once, and then a user could experiment by adding different context information interactively. Removing the retraining step also allows FS-SINR to produce estimated ranges in a fraction of the compute time compared to other approaches. Compared to the publicly released implementation of LE-SINR on the same hardware, FS-SINR generates range estimates from one context location and text for all species in the IUCN and S&T datasets in 2% of the time on CPU, and 6% of the time on GPU.
@@
We present qualitative results for three different species in Figure 7 where we visualize FS-SINR’s predictions as we change the number of context locations. Given only a single context location, the model does a sensible job of localizing the species on Earth. This supports the findings from Figure 3 where we observe strong performance even when only one context location is available. When more information is provided, the predicted range more closely resembles the expert-derived range shown in the first row. However, we do note that the model can still make mistakes in our low data setting, such as the erroneous predictions for the ‘Black and White Warbler’ in South America. In Figure 4 we illustrate some examples of how text information, when paired with one single context location, can influence the model predictions. We observe dramatically different predicted ranges when the text prompt encourages the model to focus on different habitat types. We note that each of the predicted ranges is still consistent with the location of the single context location provided. Finally, in Figure 6 we compare FS-SINR range predictions to other approaches, namely SINR, LS-SINR, and Active SINR. We see that for this species FS-SINR more closely resembles the expert range when only three context locations are provided. Additional qualitative examples are provided in Appendix E.
@@ 4.3 Zero-shot Evaluation
In addition to being able to generate range predictions in the few-shot setting when limited location observations are provided, FS-SINR can also make predictions when no location information is provided but only additional metadata such as an image or text describing a previously unseen species is given, i.e., the zero-shot setting. These zero-shot results are presented in Table 1 for both the IUCN and S&T datasets.
Table 1: Zero-shot results. We report zero-shot performance where no location information is provided to each model, only additional metadata, comparing to SINR (Cole et al., 2023) and LE-SINR (Hamilton et al., 2024). We denote additional metadata used by models as RT for ‘Range Text’, HT for ‘Habitat Text’, and ‘I’ for ‘Image’. TST represents ‘Test Species in Train’, indicating that a model uses location observations for the evaluation species at training time (e.g., SINR which provides an upper bound on performance), unlike other models where these species are excluded. TRT models are trained using ‘Taxonomic Rank Text’ as in Sastry et al. (2023), which are also provided with the full taxonomic description from ‘class’ to ‘species’ during evaluation. Results are presented as MAP, where higher is better.
@@
Figure 5: Range predictions with a single context image as input. We can condition FS-SINR on an arbitrary input image with no context locations or text, e.g., a held-out species (top row and bottom left), famous landmarks (bottom middle), or landscape images (bottom right).
@@
Figure 6: Qualitative comparison of range predictions for different methods. (Top) Predicted ranges from a single context location denoted as ‘◦’ and no additional metadata for the Gabar Goshawk. (Bottom) Predicted ranges from three context locations for the Black-naped Monarch. From left to right, FS-SINR (ours) with expert range inset, SINR, LE-SINR, and Active SINR. Please zoom in to see details.
@@
We report results for several variants of FS-SINR where different types of metadata are used. As a baseline, we also present the performance of SINR (row 1) where the evaluation species are part of its training set i.e., not zero-shot. We can also add data from these species to the training set of our approach which unsurprisingly boosts performance (i.e., row 3 vs. 9), though unlike SINR, FS-SINR does not have weights associated with individual species and so the impact of seeing evaluation species during training is fairly small. As a trivial baseline, we also report performance of FS-SINR (row 4) when no location or text metadata is provided, i.e., this is simply the output of the class token. As expected, this model performs poorly, but interestingly it seems to have learned some spatial prior that results in non-trivial predictions on S&T which contains bird species mostly concentrated in North America. We also compare to a version of FS-SINR (row 5) where we use taxonomic text (TRT) as in LD-SDM (Sastry et al., 2023) (see Section A.2 for further details).
@@
In all instances, our FS-SINR approach outperforms LE-SINR, even when both models are provided with the same information at inference and training time (i.e., row 6 vs. 7 or row 8 vs. 9). Confirming observations from LE-SINR, we see that range text (RT) is more informative than habitat text (HT) (i.e., row 7 vs. 9). Additionally, image information provides some non-trivial signal (i.e., row 4 vs. 10), but it is not as informative as text (i.e., row 9 vs. 10), and can negatively impact performance when more informative sources are provided (i.e., row 9 vs. 11 for the harder IUCN dataset), as the model may overfit to incorrect spurious features in the image. As we can see in Figure 5 (with additional examples in Figure A24), zero-shot image predictions can be sensible, but predicting an unobserved species’ range from a single input image is ill posed. Text descriptions of range or habitat preferences are simply much more informative than a single image.
4.4 Additional Results and Ablations
In Appendix A we provide additional experimental results for FS-SINR. There we investigate uncertainty quantification to see how well calibrated the model predictions are. We also report results using a ‘distance-weighted MAP’ metric which penalizes errors more the further they are in distance away from the actual range. This metric more closely aligns human judgment of predicted range quality.
We provide a more ecologically relevant breakdown of results in Appendix B, where we find multiple potential sources of bias in our training data toward North America and Europe, and report higher evaluation performance in these regions for FS-SINR and LE-SINR. Similarly, we see that biases in the text data potentially leads to increased performance for charismatic and well-studied mammals compared to other taxonomic groups, though providing more context locations reduces this gap. FS-SINR is somewhat robust to these biases and outperforms other approaches across almost all categories. We also find that for all approaches tested, estimating very small ranges is difficult and performance varies strongly with range sizes, though FS-SINR again shows comparatively good performance.
@@ 4.4 Additional Results and Ablations
We provide additional ablation experiments for FS-SINR in Appendix D, where we evaluate the impact of different input features, location encoders, and the amount of training data used, and explore architectural modifications such as removing the final species decoder that operates on the output of the Transformer. We observe that FS-SINR is robust to these changes, justifying its design decisions.
@@
[…]
@@ 5 Limitations
While FS-SINR outperforms other zero and few-shot approaches for species range estimation, there are some limitations. First, given a set of input context locations FS-SINR is deterministic in that it will always generate the same output range map. In practice, in the few-shot regime, the same set of points could actually be representative of many different possible range maps. An obvious extension of our work is to introduce stochasticity into the model outputs, e.g., by treating class token output from the Transformer as a latent embedding for an additional sampling step. In Figure A25 we observe that initializing FS-SINR with different random seeds during training results in diverse range predictions across the different models, and in Appendix A we show that this can be exploited to give estimates of the uncertainty of predictions when using an ensemble of FS-SINR models. We leave further exploration of this for future work. Second, at inference time, users may wish to provide example locations indicating where a specific species has not been found, i.e., confirmed absences. Currently, our model is trained using presence-only data but could be adapted to use absence information, if available, which could be denoted via a different embedding type vector, to be learned during training alongside our existing token type embeddings. However, obtaining reliable large-scale absence data for tens of thousands of species is a challenging problem.
Finally, biodiversity data and particularly global-scale citizen science datasets like the one we use to train FS-SINR can contain large biases (Geldmann et al., 2016; Hughes et al., 2021), e.g., location, temporal, and taxonomic biases, among others. We do not explicitly account for these biases during training, though we make some attempt to evaluate the impact of them in Appendix B, and thus we would caution the use of the predictions of our model in any applications that would use our range predictions in the context of biodiversity assessments. However, we note that we outperform existing and recent state-of-the-art range estimation methods, especially in the low observation data setting, and do not require any retraining at inference time.